# Dirichlet energy and focusing NLS condensates of minimal intensity

### v2

@@ 1 Introduction
In the next Subsection 1.1 we provide a detailed description of the results of the paper, while in the remaining subsections 1.2 of this introduction we discuss the connection and motivation for studying the problem coming from the theory of fNLS soliton condensates.
@@ 1.1 Detailed description of the results
In order to formulate the main theorem we need to define the notion of connectivity associated with a poly-continuum. Given a finite anchor set $E\subset{{\mathbb{H}}}$ we define the family ${\mathbb{K}}_{E}$ consisting of all poly-continua $\mathcal{K}$, where each component contains at least two different anchor points or connects an anchor $e\in E$ with a point in $\mathbb{R}$}. We denote the components as $\mathcal{K}=\sqcup_{\ell=0}^{k}\mathcal{K}_{\ell}$, where $\mathcal{K}_{\ell}$, $\ell=1,\dots,k,$ are the connected components of $\mathcal{K}$ not meeting $\mathbb{R}$}, and the notation $\mathcal{K}_{0}$ is reserved for the component of $\mathcal{K}\cup{\mathbb{R}}$ containing $\mathbb{R}$}. This partition of a poly-continuum $\mathcal{K}\in{\mathbb{K}}_{E}$ allow us to define the connectivity of $\mathcal{K}$ as an $(N+1)$ by $(N+1)$ symmetric matrix $M=M(\mathcal{K})$ (connectivity matrix) as follows: the entry $M_{i,j}=1$ means that $e_{i},e_{j}$ belong to the same component $\mathcal{K}_{\ell}$, $0\leq\ell\leq k$,   $1\leq i,j\leq N$, and $M_{i,0}=1$ means that $e_{i}$ belongs to $\mathcal{K}_{0}$; the remaining entries of $M$ are zeros. We say that poly-continua $\mathcal{K}^{(1)},\mathcal{K}^{(2)}\in{\mathbb{K}}_{E}$ have the same connectivity if $M(\mathcal{K}^{(1)})=M(\mathcal{K}^{(2)})$. We say that $\mathcal{K}^{(2)}$ has a connectivity that is (weakly) greater than (or exceeds the connectivity of) $\mathcal{K}^{(1)}$, if $M(\mathcal{K}^{(1)})_{i,j}\leq M(\mathcal{K}^{(2)})_{i,j}$, $i,j=0,\dots,N$ and we write $M(\mathcal{K}^{(1)})\preceq M(\mathcal{K}^{(2)})$; this provides a partial order in the set of allowed connectivities. It is clear that there are finitely many different connectivities within the class ${\mathbb{K}}_{E}$. For a given anchor set $E$ we define subclasses ${\mathbb{K}}_{E,M}\subset{\mathbb{K}}_{E}$, where each subclass consists of poly-continua with larger connectivity than the one defined by the matrix $M$. We can now formulate our main results.
@@
The class $\mathbb{K}_{E}$ and each of the sub-classes $\mathbb{K}_{E,M}$ with preassigned minimal connectivity are all closed in Hausdorff topology. We are interested in the issue of existence (and possibly uniqueness) of the minimizer of $\mathcal{I}$ on these subclasses: if they were compact (alas, they are not), then the existence would be a triviality given the established continuity. Thus we can formulate the first result as follows:
@@
For every connectivity $M$ the Dirichlet energy functional $\mathcal{I}:\mathbb{K}_{E,M}\to{\mathbb{R}}_{+}$ attains a minimum value at a poly-continuum $\mathfrak{F}\in\mathbb{K}_{E,M}$.
@@
with $P(z)$ a polynomial with real coefficients and such that $Q$ additionally satisfies the Boutroux condition: this means that all contour integrals of $w=\sqrt{Q(z)}$ along closed loops $\gamma$ on the Riemann surface of $w^{2}=Q(z)$ (hyperelliptic) are purely real, i.e.
@@
is well–defined (i.e. independent of the path of integration), harmonic away from its zero level set and behaves near $z=\infty$ like
@@
Any minimizer, $\mathfrak{F}$, of the Dirichlet energy within a class $\mathbb{K}_{E,M}$ consists of the zero level curves of the function $V(z)$ (1.8) associated with a suitable Boutroux quadratic differential, $Q$, of quasi-momentum type (1.6). Furthermore the minimum energy $\mathcal{I}(\mathfrak{F}_{Q})$ equals the coefficient $I_{Q}$ in the expansion (1.9).
@@
The final result consists in answering in a partial way the issue of uniqueness. For fixed set of anchors $E$ we choose one of the finitely many Boutroux quadratic differential of quasi-momentum type and $\mathfrak{F}_{Q}=V^{-1}(\{0\})$, see (1.10). This poly-continuum defines a certain connectivity class $M(\mathfrak{F}_{Q})$, where we recall that $M(\mathfrak{F}_{Q})$ denotes the connectivity matrix associated with the poly-continuum $\mathfrak{F}_{Q}$, see previous section. Then
@@
In the class $\mathbb{K}_{E,M({\mathfrak{F}_{Q}})}$ the minimum is unique and is attained at $\mathfrak{F}_{Q}$.
@@ 1.3 Motivation: soliton condensates for integrable equations
The idea of soliton gas for Korteweg de Vries (KdV) equation goes back to 1971 paper [40] of V. Zakharov, where he calculated the effective velocity of a trial soliton propagating on a multi-soliton background. This background modifies the average speed of the (free) trial soliton because of its repeated interactions with the background solitons, each of which can be regarded as an instantaneous shift of the center (aka as phase shift) of the trial soliton. In the modern language, the setting of [40] corresponds to a diluted KdV soliton gas. In order to study a dense KdV soliton gas, a novel approach was suggested by G. El in [5]. This approach is based on styding the finite gap solutions for the KdV, defined by some hyperelliptic Riemann surface $\mathfrak{R}$, where the number $N$ of the bands is growing but the size of the (bounded) bands simultaneously go to zero at a certain exponential (in $N$) rate. Each individual decaying band correspond to a soliton in this limit, but the key thing is the right scaling of the decaying bands, which could be found in an earlier work [38] of S. Venakides. Such a limit is called the thermodynamic limit in [5]. Among the main results of [5] are the so called Nonlinear Dispersion Relations (NDR), which define the continualized limits $u(z):\Gamma^{+}\to\bar{\mathbb{R}}^{+}$ and $v(z):\Gamma^{+}\to\bar{\mathbb{R}}$ of scaled wavenumbers and of scaled frequences respectively of the finite gap KdV solutions in the thermodynamic limit. In spectral theory of soliton gases, $u(z)$ is called the density of states (DOS) and $v(z)$ - the density of fluxes (DOF).
We note that the spectral problem for the KdV is self adjoint so that all the spectral bands of $\mathfrak{R}$ are on $\mathbb{R}$}. Deriving the NDR for a non self adjoint problem, such as, for example, Zakharov-Shabat problem for the fNLS was achieved in [7]. Since the spectral bands of $\mathfrak{R}$ are now in $\mathbb{C}$} (due to Schwarz symmetry, we could restrict our attention to the upper half plane $\mathbb{H}$}} only), the (general) NDR for fNLS soliton gases are complex. For example, the general first NDR for the fNLS soliton gas with a one dimensional accumulation set (an arc) $\Gamma^{+}\subset{{\mathbb{H}}}$ is
@@
To have another look on the NDR (1.15), (1.17), we remind that the wavenumbers and frequences of finite gap solutions are represented by periods of quasimomentum ${\mathrm{d}}p_{g}$ and quasienergy ${\mathrm{d}}q_{g}$ meromorphic differentials on $\mathfrak{R}$ of genus $g$ that are normalized in such a way that all their periods are real (real normalized or Boutroux differentials). This normalization uniquely defines ${\mathrm{d}}p_{g},{\mathrm{d}}q_{g}$ (their principlal parts at singular points are fixed). If $A$ cycles are properly oriented small loops around the corresponding shrinking bands of $\mathfrak{R}$, then the $A$ and $B$ periods of ${\mathrm{d}}p_{g}$ are called the solitonic and the carrier wavenumbers of the corresponding finite gap solutions, see [7]. Then the general first NDR (1.15) can be viewed as the thermodynamic limit of the Riemann Bilinear relations between ${\mathrm{d}}p_{g}$ and the normalized holomorphic differentials of $\mathfrak{R}$, see [37], [19]. Similarly, one can obtain (1.17) from ${\mathrm{d}}q_{g}$. One can also observe that the kernel of the integral operator in (1.15)-(1.17) was obtained as the thermodynamic limit of the Riemann Period matrix of $\mathfrak{R}$ of the genus $g$ as $g\to\infty$ (subject to some technical restrictions), where the centers of the corresponding bands of $\mathfrak{R}$ approach the values of $w,z\in\Gamma^{+}$ respectively ([36]).
@@
where the minimum is taken among all $\sigma\in C(\Gamma^{+}),\sigma\geq 0$, and all nonnegative Borel measures on $\Gamma^{+}$. Since for the minimizer $\mu_{\sigma}$ of $J_{\sigma}$ we have $J_{\sigma}[d\mu_{\sigma}]=\int_{\Gamma^{+}}\mathrm{Im}\,(z)\,d\mu_{\sigma}(z)$, and the latter expression is propotional to the average intentensity of the condensate ([36]), we conclude that the soliton condensate has the maximal average intensity among all soliton gases with spectral support $\Gamma^{+}$. That observation naturally led to the following question that triggered the work on this paper. Let a compact $\Gamma^{+}\subset{{\mathbb{H}}}$ be a finite collection of piece-wise smooth contours with the fixed endpoints comprising the set $E\subset{{\mathbb{H}}}$. For a given $E$, find $\Gamma^{+}\subset{{\mathbb{H}}}$ that maximizes the Green energy of the soliton condensate defined by $\Gamma^{+}\subset{{\mathbb{H}}}$ with the external field $2\mathrm{Im}\,z$. Equivalently, one can ask of $\Gamma^{+}\subset{{\mathbb{H}}}$ that minimizes the average intensity $\mathbb{I}=\mathbb{I}(\Gamma^{+})$) of the fNLS soliton condensate, defined by $\Gamma^{+}$. As it was discussed above, this problem can be considered as some generalized version of the Chebotarev’s continuum problem.
@@ Acknowledgements.
The authors would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, UK, and Northumbria University, Newcastle, UK, for support and hospitality during the satelite programme “Emergent phenomena in nonlinear dispersive waves”, where work on this paper was undertaken. AT was supported by a Simons fellowship for his visit. The authors also gratefully thank Evguenii Rakhmanov and Arno Kuijlaars for helpful discussions. The work of MB was supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) grant RGPIN-2023-04747. The work of AT was supported in part by the NSF Grants DMS-2009647, DMS-2407080 and by Simons Foundation award. During the revision of the work MB was supported by the Royal Society Wolfson Visiting Fellowship at the School of Mathematics, Bristol University.
@@ 2 Zakharov–Shabat equation and spectra
If $\psi(x,0)$ is a finite gap solution [1] then $\Psi(x,0;z)$ can be written explicitly in terms of Riemann Theta functions associated to a hyperelliptic Riemann surface $\mathfrak{R}_{L}$ of finite genus L1L-1 branched at $2L$ points $\{b_{1},\dots b_{L},\overline{b}_{1},\dots,\overline{b}_{L}\},\ b_{j}\in{{\mathbb{H}}}\setminus{\mathbb{R}}$.
@@
The quasienergy differential ${\mathrm{d}}\mathbf{q}$ appearing in (2.4) is also a second kind real normalized meromorphic differentials with behaviour $\left(-4z+\mathcal{O}(z^{-2})\right){\mathrm{d}}z$ as $z\to\infty$ on the main sheet of $\mathfrak{R}_{L}$, see [9].

### v3

@@ 1 Introduction
In the next Subsection 1.1 we provide a detailed description of the results of the paper, while in the remaining Subsections 1.2, 1.3 of this introduction we discuss the connection and motivation for studying the problem coming from the theory of fNLS soliton condensates.
@@ 1.1 Detailed description of the results
In order to formulate the main theorem we need to define the notion of connectivity associated with a poly-continuum. Given a finite anchor set $E\subset{{\mathbb{H}}}$ we define the family ${\mathbb{K}}_{E}$ consisting of all poly-continua $\mathcal{K}$, where each component contains at least two different anchor points or connects an anchor $e\in E$ with a point in $\mathbb{R}$}. We denote the components as $\mathcal{K}=\sqcup_{\ell=0}^{k}\mathcal{K}_{\ell}$, where $\mathcal{K}_{\ell}$, $\ell=1,\dots,k,$ are the connected components of $\mathcal{K}$ not meeting $\mathbb{R}$}, and the notation $\mathcal{K}_{0}$ is reserved for the component of $\mathcal{K}\cup{\mathbb{R}}$ containing $\mathbb{R}$}. This partition of a poly-continuum $\mathcal{K}\in{\mathbb{K}}_{E}$ allow us to define the connectivity of $\mathcal{K}$ as an $(N+1)$ by $(N+1)$ symmetric matrix $M=M(\mathcal{K})$ (connectivity matrix) below.
Definition 1.1
  • Let $\mathcal{K}$ be a poly-continuum in $\overline{{\mathbb{H}}}$ containing $E=\{e_{1},\dots,e_{N}\}$ and $\mathcal{K}_{0}$ the connected component of $\mathcal{K}\cup{\mathbb{R}}$ that contains $\mathbb{R}$}. The connectivity matrix (or pattern) $M(\mathcal{K})$ is the $(N+1)\times(N+1)$ matrix whose $(i,j)$ entry is
[…]
    $M_{0,i}(\mathcal{K})$ $=M_{i,0}(\mathcal{K})=\left\{\begin{array}[]{cc}1&\text{$e_{i}$ belongs to $\mathcal{K}_{0}$.}\\ 0&\text{otherwise}.\end{array}\right.,\ \ i=1,\dots,N$ (1.9)
    $M_{0,0}(\mathcal{K})$ $=1.$ (1.10)
  • An admissible connectivity matrix $M$ is a symmetric square matrix of size $N+1$ with entries either $0,1$ and with at least two $1$’s in each row and each column.
  • We say that poly-continua $\mathcal{K}^{(1)},\mathcal{K}^{(2)}\in{\mathbb{K}}_{E}$ have the same connectivity if $M(\mathcal{K}^{(1)})=M(\mathcal{K}^{(2)})$. We say that $\mathcal{K}^{(2)}$ has a connectivity that is (weakly) greater than (or exceeds the connectivity of) $\mathcal{K}^{(1)}$, if $M(\mathcal{K}^{(1)})_{i,j}\leq M(\mathcal{K}^{(2)})_{i,j}$, $i,j=0,\dots,N$ and we write $M(\mathcal{K}^{(1)})\preceq M(\mathcal{K}^{(2)})$
The above provides a partial order in the set of allowed connectivities. It is clear that there are finitely many different connectivities within the class ${\mathbb{K}}_{E}$. Finally, for a fixed admissible connectivity matrix $M_{0}$ we define define subclasses ${\mathbb{K}}_{E,M_{0}}\subset{\mathbb{K}}_{E}$, where each subclass consists of poly-continua with larger connectivity than the one defined by the matrix $M_{0}$:
$\mathbb{K}_{E,M_{0}}:=\bigg\{\mathcal{K}\text{ poly-continuum in $\overline{{\mathbb{H}}}$}:\ E\subset\mathcal{K},M_{0}\preceq M(\mathcal{K})\bigg\}.$ (1.11)
We can now formulate our main results.
@@
The class $\mathbb{K}_{E}$ and each of the sub-classes $\mathbb{K}_{E,M_{0}}$ in Def. 1.11 with preassigned minimal connectivity are all closed in Hausdorff topology. We are interested in the issue of existence (and possibly uniqueness) of the minimizer of $\mathcal{I}$ on these subclasses: if they were compact (alas, they are not), then the existence would be a triviality given the established continuity. Thus we can formulate the first result as follows:
@@
For every admissible connectivity $M$ (Def. 1.1) the Dirichlet energy functional $\mathcal{I}:\mathbb{K}_{E,M}\to{\mathbb{R}}_{+}$ in (1.12) attains a minimum value at a poly-continuum $\mathfrak{F}\in\mathbb{K}_{E,M}$, see (1.11).
@@
with $P(z)$ a polynomial with real coefficients and such that $Q$ additionally satisfies the Boutroux condition: this means that all contour integrals of $w=\sqrt{Q(z)}$ along closed loops $\gamma$ on the hyperelliptic Riemann surface of $w^{2}=Q(z)$ are purely real, i.e.
@@
is well–defined (i.e. independent of the path of integration), harmonic away from its zero level set and behaves near $z=\infty$ (in the upper half plane $\mathrm{Im}\,z>0$) like
@@
Let $M$ denote an admissible connectivity pattern (Def. 1.1) and $\mathbb{K}_{E,M}$ the class defined as in (1.11). Let $\mathfrak{F}\in\mathbb{K}_{E,M}$ denote any minimizer of the Dirichlet energy $\mathcal{I}$ (1.12) within the class $\mathbb{K}_{E,M}$ (see Theorem 1.2). Then $\mathfrak{F}$ consists of the zero level curves of the function $V(z)$ (1.16) associated with a suitable Boutroux quadratic differential, $Q$, of quasi-momentum type (1.14). Furthermore the minimum energy $\mathcal{I}(\mathfrak{F}_{Q})$ equals the coefficient $I_{Q}$ in the expansion (1.17).
@@
The final result consists in answering in a partial way the issue of uniqueness. For fixed set of anchors $E$ we choose one of the finitely many Boutroux quadratic differential of quasi-momentum type and $\mathfrak{F}_{Q}=V^{-1}(\{0\})$, see (1.18). This poly-continuum defines a certain connectivity class $M(\mathfrak{F}_{Q})$, where we recall that $M(\mathfrak{F}_{Q})$ denotes the connectivity matrix in Def. 1.1. Then
@@
In the class $\mathbb{K}_{E,M({\mathfrak{F}_{Q}})}$ (1.11) the minimum is unique and is attained at $\mathfrak{F}_{Q}$.
@@ 1.3 Motivation: soliton condensates for integrable equations
The idea of soliton gas for Korteweg de Vries (KdV) equation goes back to 1971 paper [40] of V. Zakharov, where he calculated the effective velocity of a trial soliton propagating on a multi-soliton background. This background modifies the average speed of the (free) trial soliton because of its repeated interactions with the background solitons, each of which can be regarded as an instantaneous shift of the center (aka as phase shift) of the trial soliton. In the modern language, the setting of [40] corresponds to a diluted KdV soliton gas. In order to study a dense KdV soliton gas, a novel approach was suggested by G. El in [5]. This approach is based on studying the finite gap solutions for the KdV, defined by some hyperelliptic Riemann surface $\mathfrak{R}$, where the number $N$ of the bands is growing but the size of the (bounded) bands simultaneously go to zero at a certain exponential (in $N$) rate. Each individual decaying band correspond to a soliton in this limit, but the key thing is the right scaling of the decaying bands, which could be found in an earlier work [38] of S. Venakides. Such a limit is called the thermodynamic limit in [5]. Among the main results of [5] are the so called Nonlinear Dispersion Relations (NDR), which define the “continualized” limits $u(z):\Gamma^{+}\to\bar{\mathbb{R}}^{+}$ and $v(z):\Gamma^{+}\to\bar{\mathbb{R}}$ of scaled wave-numbers and of scaled frequencies respectively of the finite gap KdV solutions in the thermodynamic limit. In spectral theory of soliton gases, $u(z)$ is called the density of states (DOS) and $v(z)$ - the density of fluxes (DOF).
We note that the spectral problem for the KdV is self adjoint so that all the spectral bands of $\mathfrak{R}$ are on $\mathbb{R}$}. Deriving the NDR for a non self adjoint problem, such as, for example, Zakharov-Shabat problem for the fNLS was achieved in [7]. Since the spectral bands of $\mathfrak{R}$ are now in $\mathbb{C}$} (due to Schwartz symmetry, we could restrict our attention to the upper half plane $\mathbb{H}$}} only), the (general) NDR for fNLS soliton gases are complex. For example, the general first NDR for the fNLS soliton gas with a one dimensional accumulation set (an arc) $\Gamma^{+}\subset{{\mathbb{H}}}$ is
@@
To have another look on the NDR (1.23), (1.25), we remind that the wave-numbers and frequencies of finite gap solutions are represented by periods of quasimomentum ${\mathrm{d}}p_{g}$ and quasi-energy ${\mathrm{d}}q_{g}$ meromorphic differentials on $\mathfrak{R}$ of genus $g$ that are normalized in such a way that all their periods are real (real normalized or Boutroux differentials). This normalization uniquely defines ${\mathrm{d}}p_{g},{\mathrm{d}}q_{g}$ (their principal parts at singular points are fixed). If $A$ cycles are properly oriented small loops around the corresponding shrinking bands of $\mathfrak{R}$, then the $A$ and $B$ periods of ${\mathrm{d}}p_{g}$ are called the solitonic and the carrier wave-numbers of the corresponding finite gap solutions, see [7]. Then the general first NDR (1.23) can be viewed as the thermodynamic limit of the Riemann Bilinear relations between ${\mathrm{d}}p_{g}$ and the normalized holomorphic differentials of $\mathfrak{R}$, see [37], [19]. Similarly, one can obtain (1.25) from ${\mathrm{d}}q_{g}$. One can also observe that the kernel of the integral operator in (1.23)-(1.25) was obtained as the thermodynamic limit of the Riemann Period matrix of $\mathfrak{R}$ of the genus $g$ as $g\to\infty$ (subject to some technical restrictions), where the centers of the corresponding bands of $\mathfrak{R}$ approach the values of $w,z\in\Gamma^{+}$ respectively ([36]).
@@
where the minimum is taken among all $\sigma\in C(\Gamma^{+}),\sigma\geq 0$, and all nonnegative Borel measures on $\Gamma^{+}$. Since for the minimizer $\mu_{\sigma}$ of $J_{\sigma}$ we have $J_{\sigma}[d\mu_{\sigma}]=\int_{\Gamma^{+}}\mathrm{Im}\,(z)\,d\mu_{\sigma}(z)$, and the latter expression is proportional to the average intensity of the condensate ([36]), we conclude that the soliton condensate has the maximal average intensity among all soliton gases with spectral support $\Gamma^{+}$. That observation naturally led to the following question that triggered the work on this paper. Let a compact $\Gamma^{+}\subset{{\mathbb{H}}}$ be a finite collection of piece-wise smooth contours with the fixed endpoints comprising the set $E\subset{{\mathbb{H}}}$. For a given $E$, find $\Gamma^{+}\subset{{\mathbb{H}}}$ that maximizes the Green energy of the soliton condensate defined by $\Gamma^{+}\subset{{\mathbb{H}}}$ with the external field $2\mathrm{Im}\,z$. Equivalently, one can ask of $\Gamma^{+}\subset{{\mathbb{H}}}$ that minimizes the average intensity $\mathbb{I}=\mathbb{I}(\Gamma^{+})$) of the fNLS soliton condensate, defined by $\Gamma^{+}$. As it was discussed above, this problem can be considered as some generalized version of the Chebotarev’s continuum problem.
@@ Acknowledgements.
The authors would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, UK, and Northumbria University, Newcastle, UK, for support and hospitality during the satellite programme “Emergent phenomena in nonlinear dispersive waves”, where work on this paper was undertaken. AT was supported by a Simons fellowship for his visit. The authors also gratefully thank Evguenii Rakhmanov and Arno Kuijlaars for helpful discussions. The work of MB was supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) grant RGPIN-2023-04747. The work of AT was supported in part by the NSF Grants DMS-2009647, DMS-2407080 and by Simons Foundation award. During the revision of the work MB was supported by the Royal Society Wolfson Visiting Fellowship at the School of Mathematics, Bristol University.
@@ 2 Zakharov–Shabat equation and spectra
Note that the spectrum is independent of $t$ since, by construction, the $t$ evolution is isospectral. If $\psi(x,0)$ is a finite gap solution [1] then $\Psi(x,0;z)$ can be written explicitly in terms of Riemann Theta functions associated to a hyperelliptic Riemann surface $\mathfrak{R}_{L}$ of finite genus L1L-1 branched at $2L$ points $\{b_{1},\dots b_{L},\overline{b}_{1},\dots,\overline{b}_{L}\},\ b_{j}\in{{\mathbb{H}}}\setminus{\mathbb{R}}$.
@@
The quasi-energy differential ${\mathrm{d}}\mathbf{q}$ appearing in (2.4) is also a second kind real normalized meromorphic differentials with behaviour $\left(-4z+\mathcal{O}(z^{-2})\right){\mathrm{d}}z$ as $z\to\infty$ on the main sheet of $\mathfrak{R}_{L}$, see [9].